OpenStudy (anonymous):

How many exact second's in One full day are there if there are twenty four hours in a day & 60 minutes in an hour & sixty seconds in one minute?

OpenStudy (apoorvk):

24x60x60=answer

OpenStudy (anonymous):

well thanks. -_-

OpenStudy (mysesshou):

60 minutes per one hour 60 seconds per one minute \[time = 24 hr * \frac{60 min}{1 hr}*\frac{60 sec}{1min} = 24*60*60\] the units cancel, giving you seconds.

OpenStudy (anonymous):

^ now thaat was helpful, thank you very much.